B. Sketch the graph of the following functions and then find the absolute extreme values of each of the given interval.


1.f(x)=3-2/5x;[-4,0]


2.f(x)=x-3/4+x;[-3,1]



C. Let f(x)=2x+1. Determine if the Intermediate value theorem applies to f on the closed interval [-3,4]for k=1.

1
Expert's answer
2022-03-19T02:40:07-0400

Solution (a)


The sketch of the function "f(x)=3-\frac{2}{5x}" for the interval "[-4,0]" is shown below




The plot shows that the function "f(x)=3-\frac{3}{2x}" is undefined for "x=0" . Therefore, the point "x=0" is the point of discontinuity.


From the plot, we can see that "x=-4" is the point of absolute minima and the absolute minimum value is "f(-4)=3.1"


Absolute maximum value does not exist. No point of maxima.


Solution (b)


The sketch of the function "f(x)=x-\frac{3}{4+x}" for the interval "[-3,1]" is shown below





The plot shows that the function "f(x)=x-\frac{3}{4+x}" is undefined for "x" between the given interval "[-3,1]" .


From the plot, we can see that


"x=-3" is the point of absolute minima and the absolute minimum value is "f(-3)=-6"


"x=1" is the point of absolute maxima, and the absolute maximum value is "f(1)=0.4"



Solution (c)


Given that "f(x)=2x+1" for the interval "[-3,4]"


Here


"f(-3)=2(-3)+1\\f(-3)=-5" and "f(4)=2(4)+1\\f(4)=9"


Since


"f(-3)<0" and "f(4)>0"


Therefore, the intermediate value theorem applies to f on the closed interval "[-3,4]"